The distance required for a car to stop is directly proportiional to the square of its velocity. If a car can stop in 112.5 meters at 15 kilometers per hour, how many meters are needed to stop at 25 kilometers per hour?
Answer Choices:
A 250.75
B 298.00
C 312.50
D 337.50

Answer by Nate(3500)

distance required for stopping = d
velocity = v
~~~>
(d)/(v^2) = d / v^2
(112.5) / (15)^2 = d / v^2
112.5 / 225 = d / 625
312.5 = d
(C)